# A Survey of LLM-Driven AI Agent Communication: Protocols, Security Risks, and Defense Countermeasures

### v3

@@ Abstract
In recent years, Large-Language-Model-driven AI agents have exhibited unprecedented intelligence and adaptability, and are rapidly changing human production and life. Nowadays, agents are undergoing a new round of evolution. They no longer act as an isolated island like LLMs. Instead, they start to communicate with diverse external entities, such as other agents and tools, to perform more complex tasks collectively. Under this trend, agent communication is regarded as a foundational pillar of the future AI ecosystem, and many organizations have intensively begun to design related communication protocols (e.g., Anthropic’s MCP and Google’s A2A) within the recent few months. However, this new field exposes significant security hazards, which can cause severe damage to real-world scenarios. To help researchers quickly figure out this promising topic and benefit the future agent communication development, this paper presents a comprehensive survey of agent communication security. More precisely, we first present a clear definition of agent communication and categorize the entire lifecycle of agent communication into three stages: user-agent interaction, agent-agent communication, and agent-environment communication. Next, for each communication phase, we dissect related protocols and analyze the security risks according to the communication characteristics. Then, we summarize and outlook on the possible defense countermeasures for each risk. In addition, we conduct experiments using MCP and A2A to help readers better understand the novel vulnerabilities brought by agent communication. Finally, we discuss open issues and future directions in this promising research field.
Index Terms:
large language model, agent communication, attack, and securityI Introduction
The emergence of Large Language Models (LLMs) has led to revolutionary advancements in Artificial Intelligence (AI), exhibiting unprecedented capabilities in understanding complex tasks [327]. More importantly, LLMs greatly boosted the ideal form of AI that human expects: agents111In this paper, all agents refer to LLM-driven AI agents.. Different from LLMs that are mainly like chatbots, agents possess more comprehensive capabilities (e.g., perception, interaction, reasoning, and execution), enabling them to independently complete a real-world task. For example, when users seek to make a travel plan, LLMs can only provide the best plan in text, while agents can realize the best plan in action, such as checking the weather, buying tickets, and booking hotels. Agents greatly speed up the progress of the intelligence transformations of enterprises. Their market size is expected to increase by 46% per year [235]. It can be foreseen that agents will subvert the production and living patterns of modern society, greatly changing the future business landscape. As a result, developing and promoting agents have become a strategic planning of major powers and influential companies.
@@
Now, agents are evolving towards the direction of domain-specific entities, i.e., being customized for specific scenarios and tasks. In this context, as shown in Figure 1, a task usually requires the collaboration of multiple agents, which may be located globally on the Internet. Under this condition, agent communication becomes the foundation of the future AI ecosystem. It enables agents to find other agents with specific capabilities, access external knowledge, assign tasks, and engage in other interactions. Based on the vast market of agent communication, an increasing number of communities and companies are seizing the opportunity to contribute to the development of agent communication. In November 2024, Anthropic proposed Model Context Protocol (MCP) [19], a universal protocol that allows agents to invoke external environments, such as datasets, tools, and APIs. MCP quickly gained a great deal of attention in the recent few months. Up to now, hundreds of enterprises have announced their access to MCP, such as OpenAI [214], Google [96], Microsoft [57], Amazon [25], Alibaba [10], and Tencent [267], and MCP’s package receives over 3 million weekly downloads [21]. In April 2025, Google proposed Agent-to-Agent Protocol (A2A) [89], which enables seamless communication and collaboration among agents. Since its release, A2A has received extensive support from many enterprises, such as Microsoft [200], Atlassian [160], and PayPal [242]. It can be seen that the breakthroughs in agent communication are bringing rapid and profound changes and will become an indispensable part of the AI ecosystem.
@@
However, the rapid development of agent communication also introduces complex security risks that could cause severe damage to the AI ecosystem. For example, the collaboration of cross-organization agents significantly enlarges the attack surface, causing severe security risks, including but not limited to privacy leakage, agent spoofing, agent bullying, and Denial of Service attacks. Since the research related to agent communication is still in the nascent stage, it urgently needs a systematic review of the security problems existing in the complete agent communication lifecycle. Following this trend, this paper aims to provide a comprehensive survey of existing agent communication techniques, analyze their security risks, and discuss possible defense countermeasures. We believe this work can help a broad range of readers, such as researchers who are devoted to agent development and beginners who have just started their journey in AI.
@@
We proposed the definition of agent communication for the first time, and classified it into three stages based on the communication object, which covers the entire lifecycle of agent communication.
We comprehensively studied and classified the existing agent communication protocols. Besides, we thoroughly analyzed and categorized the related security risks for each communication stage, and detailedly discussed the targeted defense countermeasures.
We conduct experiments using MCP and A2A to help readers better understand the new attack surfaces brought by agent communication. The results show that attackers can easily cause severe damage to users’ systems and privacy.
@@ II Related Work
| [170] | 2024 | Personal Agent | 4 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✗ |
| [77] | 2024 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✓ – | ✔ | ✗ | ✓ – | ✗ |
| [163] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✓ – | ✗ | ✔ | ✓ – | ✗ | ✗ |
| [143] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✔ | ✓ – | ✓ – | ✔ | ✓ – | ✗ |
| [56] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✓ – | ✓ – | ✔ | ✓ – | ✓ – | ✗ |
| [108] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✗ |
| [295] | 2025 | General IoA | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ |
| [103] | 2024 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✓ – | ✔ | ✗ | ✓ – | ✗ |
| [317] | 2025 | Agent Comm. | 6 | ✗ | ✗ | ✔ | ✔ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✗ |
| [296] | 2025 | Agent Secu. | 6 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✔ | ✔ | ✓ – | ✔ | ✔ | ✗ |
| [281] | 2025 | General Agent | 6 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✗ |
| [297] | 2024 | Agent Secu. | 6 | ✔ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✓ – | ✔ | ✗ | ✓ – | ✗ |
| [75] | 2025 | General Agent | 6 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✗ | ✗ | ✗ |
| [241] | 2025 | Agent Comm. | 7 | ✓ – | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✗ | ✗ |
| [325] | 2025 | Agent Secu. | 7 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✗ |
| [95] | 2025 | Agent Secu. | 7 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✗ | ✔ | ✗ | ✗ |
| [318] | 2025 | Agent Comm. | 8 | ✔ | ✗ | ✔ | ✔ | ✓ – | ✓ – | ✓ – | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ |
| [147] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✔ |
| [252] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ |
| [233] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ |
| [67] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ |
| [111] | 2025 | MCP Security | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✔ | ✗ | ✗ | ✔ | ✗ |
@@ II-A Overview of Novelties in This Survey
Table I summarizes the characteristics of the most relevant surveys and the differences between this survey and previous surveys. In summary, this surveys exhibit the following novelties:
This survey presents a comprehensive illustration of agent communication. Specifically, it explains why the current agent ecosystem needs communication (i.e., the predicament faced by a single agent, Section IV-A), gives the definition of agent communication (Section IV-B), and proposes a novel classification principle based on communication entity, which can cover the entire lifecyle of agent communication (Section IV-C). As a result, future studies can be included and categorized according to our survey seamlessly.
This survey exhibits a comprehensive illustration of the existing protocols related to different agent communication stages (Sections V-A, VI-A, and VII-A), including newly proposed and previously neglected protocols that have not been discussed by other surveys. Besides, we categorize these protocols based on their architecture and summarize corresponding characteristics, rather than mechanically listing each protocol. This organization method can allow any researchers interested in this field to quickly establish a preliminary but comprehensive understanding of agent communication.
This survey makes an in-depth analysis of the discovered attacks and potential risks that have not been revealed for each agent communication stage. We discuss both the risks from compromised agents and those from malicious communication objects (i.e., users and environments), fully covering the entire agent communication lifecycle. Then, we thoroughly outline the possible defense countermeasures (Sections V-D, VI-C, and VII-D) that can make future agent communication more secure.
@@ II-B Selection Principles of the Most Relevant Surveys
Principle #1: LLM-driven agents. The first and the most important is that the research object of a survey must be LLM-driven agents. This principle must be satisfied. This is because there have been many studies about multi-agent systems (MAS) before the emergence of LLMs. These agents have completely different cores and characteristics from LLM-driven agents, so discussing them benefits very little to this survey. Besides, surveys focusing on only LLMs instead of LLM-driven agents are also not listed in Table I (but we will draw on their valuable insights in other sections). This is because LLMs show significant differences from agents, for which we make a detailed illustration in Section III-C. As a result, researching LLM-driven agents is the most important principle.
Principle #2: agent communication. The second principle is that a survey should focus on or partially discuss agent communication, especially including some typical agent communication protocols such as MCP. This is because agent communication is very different from agent. However, if a survey satisfies the other two principles (i.e., LLM-driven agents and security), we still treat it as a relevant survey.
Principle #3: security. The final principle is that a survey should focus on or partially discuss agent-related security. This is because we believe that the security risks of agents still have meaning to the security risks of agent communication. The former is usually a subset of the latter, i.e., agent communication shows novel and more attack surfaces compared to agent.
@@ II-C Detailed Comparison with the Most Relevant Surveys
In this section, we will detailedly compare the most relevant surveys in Table I with our survey.
@@
Survey [170] focuses on personal agents that deeply integrate personal data and devices, exploring their potential as the main software paradigm for future personal computers. It only partially mentions the security risks related to personal agents in a section. Besides, these risks only belong to the user-agent interaction phase. It also does not discuss agent communication and related security risks and defenses. Survey [77] focuses on agent security instead of agent communication security. It is a security-specific paper, so its discussion of security is more comprehensive compared to [170]. However, the main body of its discussion is about the interaction between user and agent (U-A), without enough consideration about agent-agent (A-A) or agent-environment (A-E) interaction, which have significantly different characteristics. Besides, it also does not include any protocols related to agent communication. Survey [163] also focuses on agent security instead of agent communication security, which is similar to [77]. This survey focuses on single-agent systems and partially discusses multi-agent collaboration. It does not consider agent communication, related protocols, and enough security analyses about A-A and A-E. Survey [143] systematically summarizes seven security challenges for multi-agent systems. As shown in Table I, its main focus is on A-A, and only partially discusses U-A and A-E, which is not comprehensive. Besides, it does not consider agent communication and related protocols. Survey [56] proposes four knowledge gaps faced by agents, which mainly fall within U-A, partially discussing A-A and A-E. Besides, it does not consider agent communication or any related protocols. The defense prospect is also limited. Survey [108] focuses on the security risks of U-A and A-E, such as malicious API. It does not consider agent communication and related protocols. Besides, its security analyses are also not comprehensive enough. Survey [295] focuses on the fundamentals, applications, and challenges of IoA. Since its focus is different, agent communication and related security are only partially mentioned. Specifically, it only introduces a few related protocols and briefly analyzes related security. Besides, it also lacks the related illustration (such as definition and classification) of agent communication. Survey [103] also concentrates on the security of U-A, partially discussing A-E. It does not mention agent communication and related protocols, as well as the risks of A-A. Survey [317] focuses on agent communication architecture, which is a study of agent interaction mechanisms from a high-level and abstract view. Besides, it only partially mentioned related security and did not discuss any communication protocols. Survey [296] focuses on the security of IoA. It mentioned a few agent communication protocols (i.e., MCP, A2A, ANP, and Agora), neglecting many other important protocols. Besides, it lacks the motivation, definition, and classification of agent communication, and also does not classify protocols. According to our analyses, the security analyses (especially for U-A) are also not comprehensive enough. Survey [281] proposes the concept of “full stack safety” of agents, providing comprehensive analyses of data preparation, pre-training, post-training, deployment, and commercialization. It does not focus on agent communication security. As a result, this survey did not give a clear illustration of agent communication, only mentioned a few protocols (i.e., MCP, A2A, ANP, and Agora), and partially discussed related threats and countermeasures. Survey [297] gives a comprehensive analysis of the security of agent networks. However, it does not include the discussion of communication protocols, and lacks enough security analyses of A-A and A-E. Survey [75] does not focus on security. Instead, it concentrates on evaluating LLMs and agents. Besides, it also analyzes the architecture of some communication protocols (i.e., MCP, A2A, and ANP). We can see that it does not give a detailed illustration of agent communication, enough coverage of protocols, or a comprehensive discussion about security. Survey [241] focuses on MCP, detailedly analyzed related architectures and applications. It does not consider other communication protocols and only partially mentions security-related content. Survey [325] analyzes the threats of agents and divides them into two categories (intrinsic and extrinsic), partially covering U-A, A-A, and A-E. However, its analyses are not comprehensive enough, and it does not mention any communication protocols. Survey [95] makes a comprehensive analysis of the risks for multi-agent systems. However, its focus only falls within A-A, not considering U-A, A-E, and related protocols.

### v4

@@ Abstract
In recent years, Large-Language-Model-driven AI agents have exhibited unprecedented intelligence and adaptability. Nowadays, agents are undergoing a new round of evolution. They no longer act as an isolated island like LLMs. Instead, they start to communicate with diverse external entities, such as other agents and tools, to perform complex tasks. Under this trend, agent communication is regarded as a foundational pillar of the next communication era, and many organizations have intensively begun to design related communication protocols (e.g., Anthropic’s MCP and Google’s A2A) within the past year. However, this new field exposes significant security hazards, which can cause severe damage to real-world scenarios. To help researchers quickly figure out this promising topic and benefit the future agent communication development, this paper presents a comprehensive survey of agent communication security. More precisely, we present the first clear definition of agent communication. Besides, we propose a framework that categorizes agent communication into three classes and uses a three-layered communication architecture to illustrate how each class works. Next, for each communication class, we dissect related communication protocols and analyze the security risks, illustrating which communication layer the risks arise from. Then, we provide an outlook on the possible defense countermeasures for each risk. In addition, we conduct experiments using MCP and A2A to help readers better understand the novel vulnerabilities brought by agent communication. Finally, we discuss open issues and future directions in this promising research field. We also publish a repository that maintains a list of related papers on https://github.com/theshi-1128/awesome-agent-communication-security.
I Introduction
Computer communication has been evolving in the direction of improving interaction efficiency. Now, the emergence of Large Language Model-driven AI agents111In this paper, all agents refer to LLM-driven AI agents. [yu2025table, naveed2025comprehensive, zhao2023survey] brings revolutionary changes to this field. First, LLMs greatly boosted the evolution of agents, new entities that integrate perception, interaction, reasoning, and execution capabilities to automatically complete a real-world task. For example, when users seek to make a travel plan, LLMs can only provide recommendations in text, while agents can realize the plan in action, such as checking the weather, buying tickets, and booking hotels. Second, agents exhibit obvious domain-specific characteristics, that is, an agent is good at a certain niche field. As a result, a task usually requires the collaboration of multiple agents, which may be located globally on the Internet (as shown in Figure 1). Under these conditions, agent communication is expected to become the foundation of the next communication era and the future AI ecosystem. It enables agents to find other agents with specific capabilities, access external knowledge, assign tasks, and engage in other interactions on the Internet.
@@ I Introduction
Based on the vast market of agent communication, an increasing number of communities and companies are seizing the opportunity to contribute to its development. In November 2024, Anthropic proposed Model Context Protocol (MCP) [anthropic_mcp_2024, ray2025survey], a universal protocol that allows agents to communicate with external environments, such as datasets and tools. MCP quickly gained a great deal of attention in the recent few months. Up to now, hundreds of enterprises have announced their access to MCP, such as OpenAI [openaiagentsdk], Google [googlemcp], Microsoft [microsoftmcp], Amazon [amazonmcp], Alibaba [alibabamcp], and Tencent [tencentmcp]. In April 2025, Google proposed Agent-to-Agent Protocol (A2A) [A2A], which enables seamless communication among remote agents on the Internet. Since its release, A2A has received extensive support from many enterprises, such as Microsoft [microsoftA2A], Atlassian [AtlassianA2A], and PayPal [PayPalA2A]. The market size of agents is expected to increase by 46% per year [agentmarket].
@@
However, the rapid development of agent communication also introduces complex security risks that could cause severe damage. For example, the communication of cross-organization agents significantly enlarges the attack surface, including but not limited to privacy leakage, agent spoofing, agent bullying, and Denial of Service attacks. Since the research related to agent communication is still in the nascent stage, it urgently needs a systematic review of the security problems existing in the complete agent communication lifecycle. Following this trend, this paper aims to provide a comprehensive survey of existing agent communication techniques, analyze their security risks, and discuss possible defense countermeasures. We believe this work can help a broad range of readers, such as researchers who are devoted to agent development and beginners who have just started their journey in AI.
@@
We propose the definition of agent communication for the first time. Then, we classify agent communication into three classes based on the characteristics of communication objects and propose a three-layer communication architecture for each stage, clarifying the functional boundaries for each communication part. This fills the gap where there was a lack of a structured technical framework for agent communication.
We comprehensively studied and classified the existing 19 agent communication protocols. Besides, we thoroughly analyzed and categorized the related security risks for each communication stage and layer, and detailedly discussed the targeted defense countermeasures. This provides valuable insights for real-world deployments.
We conduct experiments using MCP and A2A to help readers better understand the new attack surfaces brought by agent communication. The results show that attacking agent communication can easily cause severe damage to the real world.
@@ II Related Work
| [li2024personal] | 2024 | Personal Agent | 4 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✗ |
| [gan2411navigating] | 2024 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✓ – | ✔ | ✗ | ✓ – | ✗ |
| [anonymous2025mind] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✓ – | ✗ | ✔ | ✓ – | ✗ | ✗ |
| [ko2025seven] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✔ | ✓ – | ✓ – | ✔ | ✓ – | ✗ |
| [deng2025ai] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✓ – | ✓ – | ✔ | ✓ – | ✓ – | ✗ |
| [he2025security] | 2025 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✗ |
| [wang2025internet] | 2025 | General IoA | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ |
| [he2024emerged] | 2024 | Agent Secu. | 5 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✓ – | ✔ | ✗ | ✓ – | ✗ |
| [yan2025beyond] | 2025 | Agent Comm. | 6 | ✗ | ✗ | ✔ | ✔ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✗ |
| [wang2025security] | 2025 | Agent Secu. | 6 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✔ | ✔ | ✓ – | ✔ | ✔ | ✗ |
| [wang2025comprehensive] | 2025 | General Agent | 6 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✗ |
| [wang2025large] | 2024 | Agent Secu. | 6 | ✔ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✓ – | ✔ | ✗ | ✓ – | ✗ |
| [ferrag2025llm] | 2025 | General Agent | 6 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✗ | ✗ | ✗ |
| [sarkar2025survey] | 2025 | Agent Comm. | 7 | ✓ – | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✗ | ✗ |
| [yu2025survey] | 2025 | Agent Secu. | 7 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✓ – | ✗ |
| [hammond2025multi] | 2025 | Agent Secu. | 7 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✔ | ✗ | ✗ | ✔ | ✗ | ✗ |
| [yang2025survey] | 2025 | Agent Comm. | 8 | ✔ | ✗ | ✔ | ✔ | ✓ – | ✓ – | ✓ – | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ |
| [krishnan2025advancing] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✔ |
| [singh2025survey] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ |
| [ray2025review] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ | ✓ – | ✗ | ✗ |
| [ehtesham2025survey] | 2025 | Agent Comm. | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ | ✓ – | ✓ – | ✗ |
| [hou2025model] | 2025 | MCP Security | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✔ | ✗ | ✗ | ✔ | ✗ |
| [zhao2025mcp] | 2025 | MCP Security | 8 | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ – | ✗ | ✗ | ✔ | ✗ | ✗ | ✔ | ✔ |
@@ II-A Overview of Novelties in This Survey
Table I summarizes the characteristics of the most relevant surveys and the differences between this survey and previous surveys. In summary, this survey exhibits the following novelties:
This survey presents a comprehensive illustration of agent communication. Specifically, it defines agent communication for the first time (Section IV-A), proposes a novel classification principle based on communication objects, which can cover the entire lifecycle of agent communication (Section IV-B), and uses a three-layered communication architecture to illustrate how each communication class is supported (Section IV-D). As a result, future studies, including protocols, attacks, and defenses, can be systematically organized.
This survey exhibits a comprehensive illustration of the existing 19 protocols related to different agent communication stages (Sections V-A, VI-A, and VII-A). Besides, we further categorize these protocols based on their characteristics, rather than mechanically listing each protocol. This organization method can allow any researchers interested in this field to quickly establish a preliminary but comprehensive understanding of agent communication.
This survey makes an in-depth analysis of the discovered attacks and potential risks that have not been revealed for each agent communication stage (Sections V-B, VI-B, VII-B). We clearly categorize each risk into different communication layers, which fully cover the entire agent communication lifecycle. Then, we thoroughly outline the possible defense countermeasures (Sections V-C, VI-C, and VII-C) that can make future agent communication more secure.
@@
This survey comprehensively discusses the future direction of agent communication from perspectives of technique and law, which can provide practical benefit for the wide adoption of agent communication.
@@ II-B Selection Principles of the Most Relevant Surveys
Principle #1: LLM-driven agents. The first and the most important is that the research object of a survey must be LLM-driven agents. This principle must be satisfied. This is because there have been many studies about multi-agent systems (MAS) before the emergence of LLMs. These agents have completely different cores and characteristics from LLM-driven agents, so discussing them benefits little for this survey. Besides, surveys focusing on only LLMs instead of LLM-driven agents are also not listed in Table I (but we will draw on their valuable insights in other sections). This is because LLMs show significant differences from agents, for which we make a detailed illustration in Section III-B3. As a result, researching LLM-driven agents is the most important principle.
Principle #2: agent communication. The second principle is that a survey should focus on or partially discuss agent communication, especially including some typical agent communication protocols such as MCP. This is because agent communication is very different from agents. However, if a survey satisfies the other two principles (i.e., LLM-driven agents and security), we still treat it as a relevant survey.
Principle #3: security. The final principle is that a survey should focus on or partially discuss agent-related security. This is because we believe that the security risks of agents still have meaning to the security risks of agent communication. The former is usually a subset of the latter, i.e., agent communication shows novel and more attack surfaces compared to agents.
@@ II-C Detailed Comparison with the Most Relevant Surveys
In this section, we will compare the most relevant surveys in Table I with our survey.
@@
Survey [li2024personal] focuses on personal agents that deeply integrate personal data and devices, exploring their potential as the main software paradigm for future personal computers. It only partially mentions the security risks related to personal agents in a section. Besides, these risks only belong to the user-agent interaction phase. It also does not discuss agent communication and related security risks and defenses. Survey [gan2411navigating] focuses on agent security instead of agent communication security. It is a security-specific paper, so its discussion of security is more comprehensive compared to [li2024personal]. However, the main body of its discussion is about the interaction between user and agent (U-A), without enough consideration about agent-agent (A-A) or agent-environment (A-E) interaction, which have significantly different characteristics. Besides, it also does not include any protocols related to agent communication. Survey [anonymous2025mind] also focuses on agent security instead of agent communication security, which is similar to [gan2411navigating]. This survey focuses on single-agent systems and partially discusses multi-agent collaboration. It does not consider agent communication, related protocols, and enough security analyses about A-A and A-E. Survey [ko2025seven] systematically summarizes seven security challenges for multi-agent systems. As shown in Table I, its main focus is on A-A, and only partially discusses U-A and A-E, which is not comprehensive. Besides, it does not consider agent communication and related protocols. Survey [deng2025ai] proposes four knowledge gaps faced by agents, which mainly fall within U-A, partially discussing A-A and A-E. Besides, it does not consider agent communication or any related protocols. The defense prospect is also limited. Survey [he2025security] focuses on the security risks of U-A and A-E, such as malicious API. It does not consider agent communication and related protocols. Besides, its security analyses are also not comprehensive enough. Survey [wang2025internet] focuses on the fundamentals, applications, and challenges of IoA. Since its focus is different, agent communication and related security are only partially mentioned. Specifically, it only introduces a few related protocols and briefly analyzes related security. Besides, it also lacks the related illustration (such as definition and classification) of agent communication. Survey [he2024emerged] also concentrates on the security of U-A, partially discussing A-E. It does not mention agent communication and related protocols, as well as the risks of A-A. Survey [yan2025beyond] focuses on agent communication architecture, which is a study of agent interaction mechanisms from a high-level and abstract view. Besides, it only partially mentioned related security and did not discuss any communication protocols. Survey [wang2025security] focuses on the security of IoA. It mentioned a few agent communication protocols (i.e., MCP, A2A, ANP, and Agora), neglecting many other important protocols. Besides, it lacks the motivation, definition, and classification of agent communication, and also does not classify protocols. According to our analyses, the security analyses (especially for U-A) are also not comprehensive enough. Survey [wang2025comprehensive] proposes the concept of “full stack safety” of agents, providing comprehensive analyses of data preparation, pre-training, post-training, deployment, and commercialization. It does not focus on agent communication security. As a result, this survey did not give a clear illustration of agent communication, only mentioned a few protocols (i.e., MCP, A2A, ANP, and Agora), and partially discussed related threats and countermeasures. Survey [wang2025large] gives a comprehensive analysis of the security of agent networks. However, it does not include the discussion of communication protocols, and lacks enough security analyses of A-A and A-E. Survey [ferrag2025llm] does not focus on security. Instead, it concentrates on evaluating LLMs and agents. Besides, it also analyzes the architecture of some communication protocols (i.e., MCP, A2A, and ANP). We can see that it does not give a detailed illustration of agent communication, enough coverage of protocols, or a comprehensive discussion about security. Survey [sarkar2025survey] focuses on MCP, detailedly analyzed related architectures and applications. It does not consider other communication protocols and only partially mentions security-related content. Survey [yu2025survey] analyzes the threats of agents and divides them into two categories (intrinsic and extrinsic), partially covering U-A, A-A, and A-E. However, its analyses are not comprehensive enough, and it does not mention any communication protocols. Survey [hammond2025multi] makes a comprehensive analysis of the risks for multi-agent systems. However, its focus only falls within A-A, not considering U-A, A-E, and related protocols.